An engineering college want to create a $81,580 annual fund for course textbooks each semester during the regular school. If the fund has a 1.8% nominal interest rate compounded biannually, how much must be in the fund to start to keep it going indefinitely?
Answer #1

Effective annual rate of a nominal biannual rate = (1+2*r)1/2 -1 = (1+2*1.8%)1/2 -1 = 1.7841%
Amount of fund required to keep it going indefinitely = 81,580/1.7841% = 4,572,613.64

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well